# Improving the Efficiency of Self-Supervised Adversarial Training through Latent Clustering-Based Selection

### v1

@@ 3.2. Limitations of Existing SSAT Methods
Although SSAT methods alleviate the sample complexity barriers of adversarially robust learning (Schmidt et al., 2018) and can produce models with higher robust accuracies, they require a substantial amount of extra unlabeled data $\mathcal{S}_{u}$ to ensure effective robustness enhancement. This trend is clearly documented in the leaderboard of RobustBench (Croce et al., 2020). For example, state-of-the-art SSAT methods (Carmon et al., 2019; Gowal et al., 2021; Wang et al., 2023) either select $500$K external Tiny Imagenet data samples or generate millions of synthetic CIFAR-10-like images, both significantly larger than the original $50$K CIFAR-10 training examples. For example, the method proposed by Gowal et al. (Gowal et al., 2021) achieves around $65\%$ robust accuracy on CIFAR-10 using a WideResNet-70-16 model architecture against $\ell_{\infty}$ perturbations with $\epsilon=8/255$, increasing the robustness performance of vanilla adversarial training (Madry et al., 2017) by a large margin, but relies an extra unlabeled set of $100$M DDPM-generated data. The huge unlabeled dataset required for SSAT significantly increases memory consumption for holding the whole training dataset, which we argue is inefficient and likely to be prohibited for resource-constrained application scenarios. These observations motivate us to explore whether such a considerable amount of unlabeled data can be reduced while preserving the high model robustness attained by SSAT algorithms.
@@
Moreover, we note that SSAT requires a much longer convergence time to obtain the best-performing model, usually $2$ to $4$ times the number of training epochs compared with vanilla AT (see Figure LABEL:fig:convergence for supporting evidence). Intuitively, extracting useful and robust features through adversarial training from such a large and potentially diverse dataset is more difficult. As we will illustrate in our experiments, the slower convergence can be attributed to the additional large unlabeled datasets (with pseudo labels) that often exhibit higher variance than the original labeled data samples. Once we reduce the size of the unlabeled data involved with specifically designed techniques, faster convergence is expected for SSAT. Since vanilla adversarial training has already been criticized in prior literature (Shafahi et al., 2019; Wong et al., 2020) for its high computational costs for running multi-step PGD to solve the inner maximization problem in Equation 3, state-of-the-art SSAT algorithms incur even higher computational costs (see Table 1 for runtime comparisons), due to the slower convergence rate and larger model architecture they typically adopt. Therefore, it is essential to design efficient SSAT methods, especially for resource-constrained applications, that can better utilize the large amount of extra unlabeled data.
3.3. Problem Formulation
@@
Witnessing the data and computational inefficiency of SSAT methods, we propose to study whether the large set size of unlabeled data is inevitable for training models with good robustness performance. Inspired by the idea of coreset selection for efficient deep learning (Kaushal et al., 2019; Xia et al., 2022; Mirzasoleiman et al., 2020), we propose to strategically search for a small but essential set of unlabeled data $\mathcal{A}_{u}\subseteq\mathcal{S}_{u}$ such that self-supervised adversarial training based on $\mathcal{S}_{l}$ and the selected subset $\mathcal{A}_{u}$ can produce models with comparable robustness to those obtained using full unlabeled dataset $\mathcal{S}_{u}$. More formally, we aim to solve the following constrained optimization problem:
@@
(4) $\max_{\mathcal{A}_{u}\subseteq\mathcal{S}_{u}}\mathrm{AdvRob}_{\epsilon}\big{(}\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)\big{)},\text{ s.t. }|\mathcal{A}_{u}|\leq\alpha|\mathcal{S}_{u}|,$
@@
where $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ stands for the model learned by SSAT with $\mathcal{S}_{l}$ and $\mathcal{A}_{u}$ based on Equation 2, and $\alpha\in(0,1)$ is a predefined ratio capturing the data constraint. $\mathrm{AdvRob}_{\epsilon}(\theta)$ denotes the robustness of the model with parameters $\theta$ against $\epsilon$ perturbations:
@@
(5) $\mathrm{AdvRob}_{\epsilon}(\theta)=1-\mathbb{E}_{(\bm{x},y)\sim D_{l}}\left[\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\mathcal{L}_{0/1}(\theta,\bm{x}+\bm{\delta},y)\right]$
@@
where $\mathcal{L}_{0/1}$ denotes the 0-1 loss function. Due to the combinatorial nature and the high computation costs of AT algorithms, it is computationally hard to enumerate all the feasible subsets $\mathcal{A}_{u}$ to solve the proposed optimization problem exactly. As we will illustrate in the following sections, we design different data selection schemes with respect to the extra unlabeled dataset for SSAT that are effective in approximately solving the optimization problem in Equation 4 while only incurring negligible computational overhead.
4. Proposed Data Selection Schemes
So far, we have introduced the problem task of reducing the unlabeled dataset size to improve the efficiency of SSAT algorithms while maintaining the robustness enhancement and illustrated its importance. Motivated by the significant role of boundary-adjacent data points in optimizing model performance (Section 4.1), in this section, we design three efficient data selection schemes to address the optimization problem in Equation 4 (Sections 4.2 and 4.3).
4.1. Prioritize Boundary Unlabeled Data
@@
Achieving high model robustness with restricted data resources remains a challenging task in machine learning. It involves striking a balance among data significance, optimizing the use of labeled and unlabeled data, and tackling constraints imposed by model capacity. Inspired by the prior literature (Zhang et al., 2020; Hua et al., 2021) that emphasizes the imbalanced data importance for vanilla adversarial training (see Section 2.1 for detailed discussions of these works), we hypothesize that not all unlabeled data contribute equally to the robustness enhancement for SSAT. In particular, we propose identifying a small set of vulnerable yet valuable unlabeled data points in $\mathcal{S}_{u}$, which are close to the model’s decision boundary. Consequently, such boundary-adjacent data points are highly susceptible to label changes under small input perturbations and are inherently difficult for the model to classify. Thus, improving their classification can yield more robustness enhancement to adversarial inputs. We expect the decision boundary of the intermediate model $f_{\hat{\theta}}$ to act as an effective proxy for locating difficult-to-classify data points. Since the intermediate model is usually trained with strong standard accuracy, it is expected to preserve class semantics and provide a solid foundation for identifying these critical points. SSAT algorithms can then leverage this information to ensure these boundary-adjacent points are correctly classified, even under adversarial conditions. By focusing more on these critical boundary points with respect to the intermediate model, the final model can be trained more efficiently using SSAT algorithms while upholding robust accuracies (see Figure 6 in Appendix A for an illustration of the overall training pipeline).
@@
While a straightforward extension of existing approaches is to select unlabeled data based on how much their predictions change under perturbations found by PGD attacks, this method is computationally expensive, undermining the efficiency gains we seek. Specifically, these methods entail iterative optimization processes to pinpoint boundary points, making them less efficient, especially for large-scale datasets. Additionally, while these methods provide valuable insights, their complexity often limits their interpretability. Therefore, we seek alternatives to effectively identify boundary data points without incurring high computational overhead.
@@
1:  Input: labeled dataset $\mathcal{S}_{l}$; unlabeled dataset $\mathcal{S}_{u}$; selection ratio $\alpha$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
@@
9:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ based on Equation 2
@@
4.2. Prediction Confidence-Based Selection
@@
To identify the critical set of vulnerable data points near the model’s decision boundary while accounting for computational efficiency, scalability, and interpretability, we first propose a straightforward approach, termed as Prediction Confidence-based Selection (PCS), which utilizes the intermediate model $f_{\hat{\theta}}$ to compute a prediction confidence score for each unlabeled data point. The pseudocode for such a selection scheme is depicted in Algorithm 1. Initially, all the data points in $\mathcal{S}_{u}$ are sorted by their prediction confidence $\mathrm{Conf}(\cdot)$, with those exhibiting the lowest confidence scores being prioritized. The underlying assumption is that data points with low confidence scores are more likely to lie near the decision boundary, making them ideal candidates for our selection. Note that the parameter $\beta\in[0,1]$ is introduced to balance the ratio between boundary and non-boundary points to avoid overfitting, which is used in all our proposed selection schemes. The reason for involving such a trade-off parameter will be further discussed in Section 6.2.
@@
The biggest advantage of PCS is its high computational efficiency for ranking the unlabeled data using model confidence scores. Note that for self-supervised adversarial training, we usually have an order-wise larger collection of unlabeled data than the original labeled dataset. Thus, an efficient data selection scheme is desirable to avoid high computational overhead. Nevertheless, we discovered in our experiments that using prediction confidence may not capture the underlying structure of the data well, leading to decreased robustness enhancement when incorporated in SSAT. We hypothesize that PCS overlooks the geometric relationships and distributional properties, which are crucial for characterizing boundary-adjacent points, particularly for complex datasets. In addition, DNNs have been shown to be overconfident in their predictions (Guo et al., 2017), suggesting prediction confidence score might be an inherently biased indicator.
4.3. Latent Clustering-Based Selection
@@
To overcome the above issues, we propose latent clustering-based selection (LCS) strategies, which identify data points near the model’s decision boundary in the latent space using different clustering techniques. Our approach begins by generating latent embeddings for all unlabeled data $\bm{z}=h_{\hat{\theta}}(\bm{x})$, where $h_{\hat{\theta}}$ denotes the mapping of the input layer to the penultimate layer with respect to the intermediate model $f_{\hat{\theta}}$. The penultimate layer captures more abstract, high-level features and better represents the underlying data structure, which helps identify points near decision boundaries. It avoids the biases of overconfident predictions from the final layer, offering more reliable clustering. Here, the goal is to identify boundary-adjacent data points inferred by examining distances to cluster centroids. Points equidistant from multiple centroids are expected to be closer to decision boundaries in the latent embedding space.
@@
1:  Input: labeled dataset $\mathcal{S}_{l}$; unlabeled dataset $\mathcal{S}_{u}$; selection ratio $\alpha$; cluster number $k$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
@@
5:  $\bm{z}\leftarrow$ get latent embeddings $h_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in D_{u}$
@@
8:     $\Delta d\leftarrow$ compute $\Delta d=|d_{1}-d_{2}|$ for $\bm{x}\in D_{u}$, where $d_{1}$ and $d_{2}$ are Euclidean distances to the nearest two centroids
@@
12:     Compute posterior probabilities $\bm{p}(\bm{z})$ for each $\bm{z}$
@@
17:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ based on Equation 2
@@
LCS-KM. In this variant of LCS, we first generate latent embeddings $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{u}\}$, and then partition the $N$ unlabeled data points into $k$ clusters $\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\}$ by minimizing the within-cluster sum of squares $\sum_{j=1}^{k}\sum_{\bm{z}\in\mathcal{C}_{j}}\|\bm{z}-\bm{\mu}_{j}\|^{2}$, where $\bm{\mu}_{j}$ is the centroid of the $j$-th cluster. For each latent embedding $\bm{z}$, we compute the Euclidean distance to each cluster centroid $\|\bm{z}-\bm{\mu}_{j}\|$. Data points are selected based on the minimal difference in distance between each latent embedding $\bm{z}$ to the corresponding two closest cluster centroids $\Delta d=|d_{1}-d_{2}|$, where $d_{1}$ and $d_{2}$ are the Euclidean distances to the closest and second closest centroids, respectively. The set of unlabeled inputs with the smallest $\Delta d$ values are selected. Finally, the top $\alpha|\mathcal{S}_{u}|$ points from $\mathcal{S}_{u}$ with the smallest $\Delta d$ values form the reduced unlabeled dataset followed by the same balancing step using the ratio parameter $\beta$. As will be demonstrated in our experiments, prioritizing these strategically selected unlabeled data points during self-supervised adversarial training can achieve comparable robustness with much-improved efficiency.
@@
LCS-GMM. In this variant of LCS, we again start by computing latent embeddings for unlabeled data $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{u}\}$. Instead of using k-means clustering, we fit these latent representations using Gaussian mixture models. A GMM assumes that the data is generated from a mixture of $k$ Gaussian distributions, each with its own mean $\bm{\mu}_{j}$ and covariance matrix $\mathbf{\Sigma}_{j}$ for the $j$-th component. Mathematically, each data point $\bm{z}$ has a probability of belonging to the $j$-th Gaussian component given by the posterior probability:
$p_{j}(\bm{z})=\frac{\pi_{j}\cdot\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})}{\sum_{i=1}^{k}\pi_{i}\cdot\mathcal{N}(\bm{z}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{i})},$

### v2

@@ II-B Limitations of Prior Work
Data Inefficiency. Although incorporating unlabeled data alleviates the sample complexity barriers of adversarially robust learning [10] and can produce models with higher robust accuracies, they require a substantial amount of extra unlabeled data $\mathcal{S}_{u}$ to ensure effective robustness enhancement. Such a trend is clearly documented in the leaderboard of RobustBench [17]. For CIFAR-10 image classification tasks, state-of-the-art SSAT algorithms either select 500500K external Tiny Imagenet samples or generate millions of synthetic CIFAR-10-like images, both significantly larger than the original 5050K CIFAR-10 training images. Specifically, the method proposed by Gowal et al. [13] achieves around $65\%$ robust accuracy on CIFAR-10 using WideResNet-70-16 against $\ell_{\infty}$ perturbations with radius $\epsilon=8/255$, increasing the robustness performance of vanilla AT [7] by a large margin; however, it relies on an extra unlabeled set of 100100M DDPM-generated data. Training SSAT algorithms on the huge unlabeled dataset demands significantly more computing power, such as specialized hardware, data parallelism, or distributed training, which in turn induces much higher energy consumption and carbon footprint. In addition, handling a larger dataset in ML often means increased GPU processing load and memory usage.
@@
High Computation Cost. Apart from the data inefficiency issue, SSAT also requires a longer convergence time to obtain the best-performing model, usually $2$ to $4$ times the number of total training epochs compared with vanilla AT (see Figure 1 for supporting evidence). Intuitively, learning meaningful and robust representations through adversarial training from such a large and potentially diverse unlabeled dataset is more difficult. As our experiments illustrate, the slower convergence of SSAT can be attributed mainly to the large set of unlabeled data (paired with pseudo labels), which often induces higher gradient variance than the original labeled data samples. Once we reduce the size of the unlabeled data while retaining the essential information, SSAT algorithms are observed to converge much faster. We note that vanilla AT has already been criticized for its high computation [18, 19], due to the iterative PGD steps involved in dealing with the inner maximization problem (Equation 3). SSAT methods incur even higher computations, since they converge more slowly and usually adopt larger model architectures (see Table I for concrete runtime comparisons). Therefore, developing more efficient methods, especially for resource-constrained scenarios, becomes critical for SSAT to effectively learn the core information in the unlabeled dataset.
II-C Problem Formulation
Witnessing the data and computational inefficiency of SSAT methods, we propose to study whether the large set size of unlabeled data is inevitable for achieving high model robustness. To be more specific, we introduce two general methodology frameworks for improving the efficiency of SSAT algorithms, corresponding to the following optimization problems:
@@ II-C Problem Formulation
Strategic Selection. Inspired by the idea of coreset selection for data-efficient deep learning [20, 21, 22], the first approach aims to strategically search for a small but essential set of unlabeled data $\mathcal{A}_{u}\subseteq\mathcal{S}_{u}$ such that SSAT with the selected subset $\mathcal{A}_{u}$ can produce models with comparable robustness to those obtained using the full unlabeled dataset $\mathcal{S}_{u}$. Formally, we aim to solve the following constrained optimization problem:
@@
$\max_{\mathcal{A}_{u}\subseteq\mathcal{S}_{u}}\mathrm{AdvRob}_{\epsilon}\big(\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)\big),\text{ s.t. }|\mathcal{A}_{u}|\leq\alpha|\mathcal{S}_{u}|,$ (4)
@@
where $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$ denotes the model learned by SSAT with $\mathcal{S}_{l}$ and $\mathcal{A}_{u}$ based on Equation II-A, and $\alpha\in(0,1)$ is a predefined ratio capturing the data constraint. Here, $\mathrm{AdvRob}_{\epsilon}(\theta)$ denotes the model robustness on $\mathcal{D}_{l}$ under $\epsilon$ perturbations:
@@
$\mathrm{AdvRob}_{\epsilon}(\theta)\!=\!1\!-\!\mathbb{E}_{(\bm{x},y)\sim D_{l}}\!\left[\max_{\bm{\delta}\in\mathcal{B}_{\epsilon}(\bm{0})}\!\ell_{0/1}(\theta,\bm{x}+\bm{\delta},y)\right]$ (5)
@@
where $\ell_{0/1}$ denotes the 0-1 loss function. Due to the combinatorial nature and the high computation costs of AT algorithms, it is computationally hard to enumerate all the feasible subsets $\mathcal{A}_{u}$ to solve the proposed optimization problem exactly. As we will illustrate in Section III-B, we design different selection schemes for the extra unlabeled dataset $\mathcal{S}_{u}$ that are effective in approximately solving the optimization problem in Equation 4 while only incurring negligible computational overhead.
Guided Diffusion. The second method proposes to leverage existing pretrained diffusion models but fine-tune them to directly generate a small set of extra data with desirable properties, ready to be incorporated into the SSAT pipeline. Similar to our first strategic selection approach, we aim to achieve comparable robustness to existing SSAT while maximizing the efficiency for the utilization of the additional data:
$\max_{\mathcal{G}_{u}}\mathrm{AdvRob}_{\epsilon}\big(\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{G}_{u},\gamma)\big),\text{ s.t. }|\mathcal{G}_{u}|\leq\alpha|\mathcal{S}_{u}|,$ (6)
where $\mathcal{G}_{u}$ denotes a small set of unlabeled data generated by a fine-tuned generative model. As will be detailed in Section III-C, we adopt a pretrained DDPM [23] as the backbone while strategically fine-tuning it with novel training losses for guided data generations. Compared with the previous selection-based method, a guided diffusion approach can avoid the inefficient pre-generation of a large amount of unlabeled data, making it particularly suitable for SSAT methods that already utilize a pre-trained diffusion model for synthetic data generation.
III Improving the Efficiency of SSAT via Latent Clustering-Based Data Reduction
So far, we have formulated the problem tasks of reducing the unlabeled dataset size to improve SSAT’s efficiency while maximizing the robustness enhancement. In this section, we first explain why boundary-adjacent data points are essential for more efficient utilization of unlabeled data (Section III-A), then we detail the proposed method designs for both strategic selection and guided generation (Sections III-B and III-C).
III-A Prioritize Boundary-Adjacent Unlabeled Data
@@
Achieving high robustness with restricted data resources remains challenging, as it involves striking a balance between data significance, optimizing the use of unlabeled data, and tackling constraints imposed by model capacity. Inspired by prior literature [16, 24] that emphasizes the unbalanced data importance for vanilla adversarial training (see Appendix A for more discussions), we hypothesize that not all unlabeled data contribute equally to the robustness enhancement for SSAT. In particular, we propose to identify or generate a small set of vulnerable yet valuable unlabeled data points, which are close to the model’s decision boundary. Consequently, such boundary-adjacent data points are highly susceptible to label changes under small input perturbations and are inherently difficult for the model to classify. Thus, improving their classification can yield more robustness enhancement to adversarial inputs. We expect the decision boundary of the intermediate model $f_{\hat{\theta}}$ to act as an effective proxy function for locating difficult-to-classify data points. Since the intermediate model is usually trained with strong standard accuracy, it is expected to preserve class semantics and provide a solid foundation for identifying these critical points. By focusing more on these critical boundary points with respect to the intermediate model, the final model can be trained more efficiently using SSAT algorithms while upholding robust accuracies.
@@
While a plausible approach is to select a set of unlabeled data based on how much their label predictions change under adversarial perturbations returned by PGD attacks, it is computationally expensive, undermining the efficiency gains we aim for. In particular, such a PGD-based method entails iterative optimization processes to pinpoint boundary points, making them impractical for large-scale unlabeled datasets. Therefore, we seek better alternatives to effectively identify boundary data points without incurring high computational overhead.
@@
1:  Input: labeled $\mathcal{S}_{l}$; unlabeled $\mathcal{S}_{u}$; selection ratio $\alpha$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
@@
9:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$
@@
III-B Strategic Selection via Latent Clustering
@@
PCS. To identify a small, critical subset of boundary-adjacent data points while accounting for computational efficiency and scalability, we first propose a straightforward approach, coined prediction confidence-based selection (PCS), which makes use of the intermediate model $f_{\hat{\theta}}$ to compute a prediction confidence score for each unlabeled data point. The pseudocode for such a selection scheme is depicted in Algorithm 1. Initially, all the data points in $\mathcal{S}_{u}$ are sorted by their prediction confidence $\mathrm{Conf}(\cdot)$, with those exhibiting the lowest confidence scores being prioritized. The underlying assumption is that data points with low confidence scores are more likely to lie near the model’s decision boundary, making them ideal candidates for our selection. Note that the parameter $\beta\in[0,1]$ is introduced to balance the ratio between boundary and non-boundary points to avoid overfitting, which is used in all our proposed selection schemes. The reason for involving such a trade-off parameter will be further discussed in Section V-C.
@@
The biggest advantage of PCS is its high computational efficiency for ranking the unlabeled data. Note that for SSAT, we usually have an order-wise larger collection of unlabeled data than the original labeled dataset. Thus, an efficient selection scheme is desirable to avoid high computational overhead. Nevertheless, our experiments showed that using prediction confidence may not capture the underlying structure of the data well, leading to decreased robustness enhancement when incorporated in SSAT. We hypothesize that PCS overlooks the geometric relationships and distributional properties, which can be crucial for characterizing boundary-adjacent points. In addition, DNNs have been shown to be overconfident in their predictions [25], suggesting that prediction confidence scores returned by neural network models might be inherently biased.
@@
LCS. To overcome the aforementioned issues, we introduce latent clustering-based selection (LCS) schemes to pinpoint desirable data in the model’s latent space using different clustering techniques. Our approach begins by generating latent embeddings for all $\bm{z}=h_{\hat{\theta}}(\bm{x})$ with $\bm{x}\in\mathcal{S}_{u}$, where $h_{\hat{\theta}}$ denotes the mapping of the input layer to the penultimate layer with respect to the intermediate model $f_{\hat{\theta}}$. The penultimate layer captures more abstract, high-level features and better represents the underlying data structure, which helps identify points near decision boundaries. It avoids the biases of overconfident predictions from the final layer, offering more reliable clustering. Here, the goal is to identify boundary-adjacent data points inferred by examining distances to cluster centroids. Points equidistant from multiple centroids are expected to be closer to decision boundaries in the latent embedding space.
@@
1:  Input: labeled $\mathcal{S}_{l}$; unlabeled $\mathcal{S}_{u}$; selection ratio $\alpha$; cluster number $k$; perturbation size $\epsilon$; boundary ratio $\beta$; per-batch ratio $\gamma$
@@
5:  $\bm{z}\leftarrow$ get latent embeddings $h_{\hat{\theta}}(\bm{x})$ for $\bm{x}\in\mathcal{S}_{u}$
@@
8:   $\Delta d\leftarrow$ compute $\Delta d=|d_{1}-d_{2}|$ for $\bm{x}\in\mathcal{S}_{u}$, where $d_{1}$ and $d_{2}$ are Euclidean distances to the nearest two centroids
@@
12:   Compute GMM posterior probability $\bm{p}$ for each $\bm{z}$
@@
17:  ${\theta}_{\mathrm{final}}\leftarrow$ $\mathrm{SSAT}(\mathcal{S}_{l},\mathcal{A}_{u},\gamma)$
@@
LCS-KM. In this variant of LCS, we first generate latent embeddings $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{u}\}$, and then partition the unlabeled dataset into $k$ clusters $\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\}$ by minimizing the within-cluster sum of squares $\sum_{j=1}^{k}\sum_{\bm{z}\in\mathcal{C}_{j}}\|\bm{z}-\bm{\mu}_{j}\|^{2}$, where $\bm{\mu}_{j}$ is the centroid of the $j$-th cluster. For each $\bm{z}$, we compute the Euclidean distance to each cluster centroid $\|\bm{z}-\bm{\mu}_{j}\|$. Data points are selected based on the minimal difference in distance between each latent embedding $\bm{z}$ to the corresponding two closest cluster centroids $\Delta d=|d_{1}-d_{2}|$, where $d_{1}$ and $d_{2}$ are the Euclidean distances to the closest and second closest centroids, respectively. Finally, the top $\alpha|\mathcal{S}_{u}|$ points from $\mathcal{S}_{u}$ with the smallest $\Delta d$ values form the reduced unlabeled dataset, followed by the same balancing step using the ratio parameter $\beta$. As will be demonstrated in our experiments, prioritizing these strategically selected unlabeled data points during semi-supervised adversarial training can achieve comparable robustness with much-improved efficiency.
@@
LCS-GMM. In this variant of LCS, we again start by computing latent embeddings for unlabeled data $\{\bm{z}=h_{\hat{\theta}}(\bm{x}):\bm{x}\in\mathcal{S}_{u}\}$. Instead of using k-means clustering, we fit these latent representations using Gaussian mixture models (GMMs). A GMM assumes that the data is generated from a mixture of $k$ Gaussian distributions, each with its own mean $\bm{\mu}_{j}$ and covariance matrix $\mathbf{\Sigma}_{j}$ for the $j$-th component. Mathematically, each data point $\bm{z}$ has a probability of belonging to the $j$-th Gaussian component given by the posterior probability:
$p_{j}(\bm{z})=\frac{\pi_{j}\cdot\mathcal{N}(\bm{z}\mid\bm{\mu}_{j},\mathbf{\Sigma}_{j})}{\sum_{i=1}^{k}\pi_{i}\cdot\mathcal{N}(\bm{z}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{i})},$ (7)